(1-1/2)(1-1/3)(1-1/4)...(1-1/9) = 1/9


may i know how this is so?

Grade:

2 Answers

AskiitiansExpert Abhinav Batra
13 years ago

Dear s adhitya

(1-1/2)(1-1/3)(1-1/4)(1-1/5).....(1-1/9)

=1/2*2/3*3/4*......8/9

=1/9(As the denominator and numerator of two consecutive terms cancel except numerator of first and denominator of last term)

 

ALL THE BEST

AskIITIANS Expert

Abhinav Batra

Parashara K.V.R
11 years ago

(1-1/2)(1-1/3)(1-1/4)......(1-1/9)

=(1/2)(2/3)(3/4)....(7/8)(8/9)

=1/9